Qs=150+14P and demand function P= 16.875-0.0625Q with 10 number of sellers and buyers.

Find equilibrium price and quantity

Find price elasticity of demand and and supply at equilibrium price and quantity




1
Expert's answer
2021-01-28T06:53:59-0500

a)Let's first rearrange our demand function:


"0.0625Q=16.875-P,""Q_d=270-16P."


Let's find the equilibrium price and quantity. In the equilibrium, "Q_d=Q_s" and we can find the equilibrium price:


"270-16P_E =150+14P_E,""30P_E=120,""P_E=\$4."

Then, we can find the equilibrium quantity by substituting "P_E" into the equation for "Q_d":


"Q_d=Q_s=Q_E=270-16\cdot4=206."


b) Let's rewrite our equations for "Q_d" and "Q_s" in slope-intercept form:


"P=16.875-0.0625Q_d,""P=0.071Q_s-10.71."

The price elasticity of demand can be found as follows:


"E_d=(\dfrac{-1}{slope})(\dfrac{P}{Q_d})=(\dfrac{-1}{(-0.0625)})(\dfrac{4}{206})=0.31"

The price elasticity of supply can be found as follows:


"E_s=(\dfrac{1}{slope})(\dfrac{P}{Q_s})=(\dfrac{1}{0.071})(\dfrac{4}{206})=0.27"